# Neural Networks for 3D Characterization of AGATA Crystals

### v1

@@ 2.4 Evaluation Using Mean Standard Deviation
    Collect all super-traces (from the validation set or full dataset inference) predicted by the NN model to be at that position. For comparison, super-traces derived from the PSCS algorithm were also analyzed.
@@
    Normalize the energy scale of all collected super-traces (scaling the maximum of the signal to 1000 units).
@@
A lower MSD indicates higher consistency among the super-traces predicted at the same position. This metric was used both to evaluate the overall quality of the NN predictions compared to PSCS and as part of an iterative filtering process to remove outlier signals before generating the final experimental bases. The filtering involved repeatedly calculating the MSD, identifying and removing the super-traces most different from the mean super-trace at that position, and stopping when the improvement in MSD plateaued (monitored via its second derivative). Similar filtering method was applied to the PSCS signals [13].
@@ 2.5 Validation using Emulated 1-D Scans
Figure 5 compares the results of the two methods (the error in this case is the mean absolute error). The error on the previously unknown axis for the 1D-trained model is only marginally higher than the error for the 2D-trained model. This suggests that the model learns the underlying relationship between pulse shape and 3D position effectively, and is capable of generalizing to predict the unknown coordinate(s) with reasonable accuracy, validating the approach of training with partial ground truth for 3D position prediction. This test provides confidence that training on two 2D scans can yield a model capable of full 3D prediction, potentially enabling future, faster 1D scanning campaigns.
Figure 5: Error on the validation dataset versus segment number. Comparison between a model trained on the standard 2D dataset (blue) and a model trained on the simulated 1D dataset (orange). The error shown is calculated only for the axis that was unknown during the 1D training but known in the original 2D dataset. The segments are labeled with a numerical index from 0 to 35. This index correlates directly with the standard AGATA nomenclature, where segments 0-5 correspond to Layer 1 (Sectors A-F), segments 6-11 correspond to Layer 2 (Sectors A-F), and so on, up to segments 30-35 for Layer 6.
@@ 2.6 Implementation Details
All code for data processing, model training, and analysis was written in Python 3. The neural networks were implemented and trained using the TensorFlow library [28] with the Keras API [29]. Training was performed on a dedicated machine learning server at IP2I, utilizing NVIDIA RTX6000 GPUs.
@@ 3 Results
The performance of the NN model is evaluated based on hit distributions, signal consistency (measured by MSD), and by comparing PSA results obtained using different signal bases. This section will primarily present the results for the A005 crystal. We will only use S001 crystal for the comparison of signal consistency (MSD) between the NN and PSCS algorithm as PSCS bases were not generated for the A005 crystal.
@@ 3.1 Hit Distributions and Error Energy Dependence
Figure 7 displays both the prediction error (blue, right axis) and the event statistics (green, left axis) as a function of the deposited energy for the A005 crystal. Compton events (300-480 keV) generally exhibit lower error than photopeak events (around 662 keV). This is consistent with previous findings [13] and attributed to the fact that photopeak events often involve multiple interactions within the same segment, leading to an averaged pulse shape whose effective position might deviate from the initial interaction point targeted by the scan. The error decreases for higher energy Compton events, possibly due to a reduced probability of random coincidences. The regions between the Compton edge (477 keV) and the photopeak, and above the photopeak, show high error and very low statistics, as single interactions in these energy ranges are physically unlikely.
@@ 3.2 Signal Consistency: Mean Standard Deviation (S001)
The MSD was calculated for all predicted positions. When averaged over the entire crystal volume, the MSD for the NN model is 7.8 a.u, a significant improvement over the 10.6 a.u value obtained for the PSCS algorithm. This global trend of superior consistency is illustrated visually in Fig 8, which compares the MSD per voxel for a representative slice of the S001 crystal at Z=30 mm. The map generated from the NN model (b) exhibits markedly higher homogeneity and fewer localized regions of high deviation (hot spots) compared to the map from the PSCS algorithm (a). This visual assessment is confirmed quantitatively for this specific slice, where the mean MSD was 7.4 a.u for the NN model versus 8.4 a.u for PSCS. These results, consistent with the trend observed throughout the entire crystal volume, indicate that the NN method produces a more uniform and reliable set of signal bases.
@@ 3.4 Pulse Shape Analysis Performance
The quality of the generated NN experimental bases (NN bases) was evaluated by using them within the standard AGATA PSA framework [30] (full grid search, 300-700 keV energy range) to predict the interaction positions of the scanned super-traces. The performance was compared against PSA using bases derived from the standard simulated ADL bases and simulated bases from AGATAGeFEM.
The position resolution (FWHM) achieved with the three different signal bases for the two known axes of each scan is summarized in Table 1. The experimental NN bases (3.0-3.1 mm) clearly demonstrate superior performance, yielding the best resolution overall. Among the simulated bases, the AGATAGeFEM bases (4.8-6.2 mm) surpassed the performance of the ADL bases(5.2-8.1 mm), a result consistent with previous study [7].
@@
Figures 9 and 10 show the PSA error maps for the A005 vertical scan across two different depths: the fourth layer and the back-most sixth layer, respectively. This comparison highlights both consistent performance trends and the impact of interaction depth. Across both layers, several key observations hold true. The direct prediction from the NN model itself (panel (a) in both figures) consistently provides the lowest apparent error and the best homogeneity. Among the PSA-based results, the experimental NN bases (panel b) yield a superior performance compared to the simulated AGATAGeFEM and ADL bases (panels c and d, respectively). The simulated bases, in particular, show degraded performance at radii $\geq$ 20 mm, suggesting potential issues in accurately simulating the core signal rise time in the outer regions of the crystal. A region of increased error is evident in the middle of the segments as it can be seen in layer 4 (Fig. 9 (b)). This error hotspot corresponds to areas where the transient signal strength in neighboring segments is minimal (see signal at R=22 mm in Fig. 1), reducing the intrinsic position sensitivity. This challenge particularly affects the PSA algorithms, whereas the direct NN model (panel (a)) proves more resilient, exhibiting a significantly smaller increase in error in this region. A direct comparison between the two figures illustrates the general trend of increasing error towards the back of the crystal. The overall error in layer 6 (Fig. 10) is visibly higher for all methods. This is attributed to two main factors: the larger segment volumes at the back reduce position sensitivity, and more importantly, the increased probability of random coincidences and scattering effects from preceding layers in the vertical scan. The performance degradation caused by these effects is significantly more pronounced in the final layer of the crystal.
Figure 9: PSA position error map for the A005 vertical scan across the fourth layer of the crystal. Comparison of the direct NN model prediction (a) and PSA results using NN bases (b), AGATAGeFEM bases (c), and ADL bases (d). Yellow color is above 8 mm.
Figure 10: PSA position error map for the A005 vertical scan across the sixth (back) layer of the crystal. The panel arrangement is identical to Fig. 9. Yellow color is above 8 mm.
@@ 3.5 PSA Performance Relative to NN Predictions
Given the potential contamination from random coincidences and mislabeled events in the scanning data, comparing PSA results directly to the scanned position might not be the most accurate way to assess the intrinsic quality of the bases. The NN model itself, trained on the large dataset, might provide a more robust reference position for each super-trace.
Therefore, to further evaluate the performance of each PSA bases, the results were compared to the positions predicted by the NN model itself. The 3D position resolution (FWHM) for each basis is summarized in Table 1.
@@
The analysis of PSA performance is significantly complicated by the presence of mislabeled events, such as random coincidences, which are particularly prevalent towards the back of the crystal in the vertical scan. This is evident when comparing the error maps relative to the scanned position (Fig. 10) with those benchmarked against the NN model’s predictions (Fig. 12). The substantial reduction in error in the latter case demonstrates that comparing PSA results to the NN model’s output effectively neutralizes the impact of these random coincidences, as both algorithms agree on the position of the outlier event, independent of the scan label.
@@
Figure 11: Error map comparing PSA predictions to the direct NN model predictions for the A005 vertical scan across the fourth layer. The error is the 2D distance between the position predicted by PSA(bases) and the NN model. Panels show PSA results using: NN bases (a), AGATAGeFEM bases (b), and ADL bases (c). Yellow color is above 8 mm.
Figure 12: Error map comparing PSA predictions to NN model predictions for the sixth (back) layer. The panel arrangement is identical to the Fig. 11. Yellow color is above 8 mm.
@@ 4 Discussion
The results presented demonstrate the successful application of neural networks, specifically LSTM-based architectures, for analyzing AGATA crystal scanning data obtained with the Strasbourg tables. The key achievement is the development of a method capable of predicting the full 3D interaction position from input super-traces, even when trained using data with only partial (2D) ground truth information available from the vertical and horizontal scans. This was enabled by a custom masked loss function and careful training strategies, including training separate models per segment and incorporating a Gaussian noise layer to mitigate overfitting.
@@
Despite the promising results, certain limitations should be acknowledged, while filtering methods were applied, the training data likely contained residual contamination from random coincidences, multiple-hit events, or other experimental artifacts. The robustness of the NN model against imperfections in the training data was therefore explicitly tested [31]. By deliberately corrupting 50% of the position labels within a segment’s dataset, it was found that the model could still achieve a prediction error close to that obtained with clean data (2.0 mm vs 1.9 mm), particularly when using the masked Euclidean distance loss function, which is inherently less sensitive to outliers than MSE. This result is significant because real experimental scanning data inevitably contains a fraction of mislabeled events. This result increases confidence that the NN model can effectively learn the correct signal-position relationships even from imperfect datasets, supporting the reliability of the generated experimental bases. This resilience mitigates concerns about data purity to some extent.
@@
Comparing the NN approach to the established PSCS algorithm, several advantages emerge. While PSCS provides a valuable algorithmic benchmark, the NN method demonstrated superior signal consistency (lower MSD) and resulted in better PSA performance. Computationally, training the NN models requires significant resources (around 10 hours on 3 GPUs for one crystal), but the subsequent inference step (predicting positions for the entire dataset or generating bases) is relatively fast (approx. 3 hours). This contrasts with PSCS, described as computationally intensive in its execution requiring 5 days [13]. Once trained, applying the NN model is straightforward and readily integrated into automated analysis workflows, potentially simplifying the process compared to the potentially complex PSCS algorithm.
@@
An additional useful outcome of applying the NN model is the generation of a large dataset where each experimental super-trace is associated with an estimated 3D interaction position. While the original scanning data only provides partial 2D labels, and simulations might not capture all experimental details, this NN-derived dataset offers comprehensive position estimates for a large portion of the recorded events. This dataset serves two practical purposes. First, it allows for a more detailed characterization of traditional PSA algorithm performance [31]. By comparing PSA results against the NN model’s 3D position estimates across the detector volume (as explored in Section 3.5), one can gain insights into PSA accuracy and limitations. Second, this dataset provides suitable training material for exploring machine learning models aimed directly at position prediction to replace the current PSA algorithm.
@@
The successful validation using simulated 1D scans opens new possibilities for future characterization campaigns. If a model can indeed be reliably trained using only one known coordinate per signal, it suggests that scanning could potentially be performed using three orthogonal 1D scans instead of two 2D scans. This could significantly reduce the time required for crystal characterization. The concept of using 1D scanning for detector characterization is not new and has been effectively demonstrated in the context of PET scanners using fan-beam collimators [32, 33, 34]. Interestingly, the analysis of that data often relied on machine learning models such as k-nearest neighbor algorithms or gradient-boosted decision trees. This choice of simpler models was well-suited to the nature of PET detector signals, which typically have a smaller size and less complex structure compared to the super-traces from AGATA.
@@ 5 Conclusions
This work successfully demonstrated the development and application of a machine learning approach, utilizing LSTM-based neural networks, to analyze experimental data from AGATA crystal scanning campaigns. By employing a masked loss function and tailored training strategies, the NN model effectively learned to predict 3D gamma-ray interaction positions from super-traces despite being trained on data with only 2D ground truth information.
@@
This study highlights the power of machine learning as a tool for processing complex detector signals and overcoming limitations inherent in traditional analysis methods or simulations. The results suggest that NN-based approaches can provide a more accurate and potentially more efficient pathway to generating high-fidelity experimental bases crucial for optimizing the performance of position-sensitive detectors like AGATA. Furthermore, the findings provide encouraging, albeit preliminary, evidence for the feasibility of potentially faster 1D scanning techniques for future detector characterization. This work serves as a foundational step toward developing advanced machine learning-based PSA algorithms capable of resolving multiple interactions within a single segment. Overall, this research underscores the growing role of machine learning in advancing data analysis capabilities in experimental nuclear physics.
@@ Acknowledgments
The authors would like to thank the the Institut Pluridisciplinaire Hubert Curien (IPHC), Strasbourg, for providing the experimental scanning data and the comparative PSCS bases.

### v2

@@ 2.4 Evaluation Using Mean Standard Deviation
    Collect all super-traces (from the validation set or full dataset inference) predicted by the NN model to be at that position. For comparison, super-traces derived from the PSCS algorithm were also analysed.
@@
    Normalise the energy scale of all collected super-traces (scaling the maximum of the signal to 1000 units).
@@
A lower MSD indicates higher consistency among the super-traces predicted at the same position. This metric was used both to evaluate the overall quality of the NN predictions compared to PSCS and as part of an iterative filtering process to remove outlier signals before generating the final experimental bases. The filtering involved repeatedly calculating the MSD, identifying and removing the super-traces most different from the mean super-trace at that position, and stopping when the improvement in MSD plateaued (monitored via its second derivative). Similar filtering method was applied to the PSCS signals [PSCS].
@@ 2.5 Validation using Emulated 1-D Scans
Figure 5 compares the results of the two methods, presenting the mean absolute error on the previously unknown axis for the 1D-trained model, the same error patterns are observed as in Fig. 3. The error for the 1D-trained model is only marginally higher than the error for the 2D-trained model. This suggests that the model learns the underlying relationship between pulse shape and 3D position effectively, and is capable of generalising to predict the unknown coordinate(s) with reasonable accuracy, validating the approach of training with partial ground truth for 3D position prediction. This test provides confidence that training on two 2D scans can yield a model capable of full 3D prediction, potentially enabling future, faster 1D scanning campaigns.
Figure 5: The mean absolute error on the validation dataset versus segment number. Comparison between a model trained on the standard 2D dataset (blue) and a model trained on the simulated 1D dataset (orange). The error shown is calculated only for the axis that was unknown during the 1D training but known in the original 2D dataset. The segment indexing follows the same nomenclature as defined for Fig. 3.
@@ 2.6 Implementation Details
All code for data processing, model training, and analysis was written in Python 3. The neural networks were implemented and trained using the TensorFlow library [abadi2016tensorflow] with the Keras API [keras2015-library]. Training was performed on a dedicated machine learning server at IP2I, utilising NVIDIA RTX6000 GPUs.
@@ 3 Results
The performance of the NN model is evaluated based on hit distributions, signal consistency (measured by MSD (see section 2.4.)), and by comparing PSA results obtained using different signal bases. This section will primarily present the results for the A005 crystal. We will only use S001 crystal for the comparison of signal consistency (MSD) between the NN and PSCS algorithm as PSCS bases were not generated for the A005 crystal.
@@ 3.1 Hit Distributions and Error Energy Dependence
Figure 7 displays both the prediction error (blue, right axis) and the event statistics (green, left axis) as a function of the deposited energy for the A005 crystal. Compton events (300-480 keV) generally exhibit lower error than photopeak events (around 662 keV). This is consistent with previous findings [PSCS] and attributed to the fact that photopeak events often involve multiple interactions within the same segment, leading to an averaged pulse shape whose effective position might deviate from the initial interaction point targeted by the scan. The error decreases for higher energy Compton events, possibly due to a reduced probability of random coincidences. The regions between the Compton edge (477 keV) and the photopeak, and above the photopeak, show high error and very low statistics, as single interactions in these energy ranges are physically unlikely.
@@ 3.2 Signal Consistency: Mean Standard Deviation (S001)
The MSD was calculated for all predicted positions. When averaged over the entire crystal volume, the MSD for the NN model is 7.8 a.u, a significant improvement over the 10.6 a.u value obtained for the PSCS algorithm. This global trend of superior consistency is illustrated visually in Fig 8, which compares the MSD per voxel for a representative slice of the S001 crystal at Z=30 mm. The map generated from the NN model (b) exhibits markedly higher homogeneity and fewer localised regions of high deviation (hot spots) compared to the map from the PSCS algorithm (a). This visual assessment is confirmed quantitatively for this specific slice, where the mean MSD was 7.4 a.u for the NN model versus 8.4 a.u for PSCS. These results, consistent with the trend observed throughout the entire crystal volume, indicate that the NN method produces a more uniform and reliable set of signal bases.
@@ 3.4 Pulse Shape Analysis Performance
The quality of the generated NN experimental bases (NN bases) was evaluated by using them within the standard AGATA PSA framework [stez] (full grid search, 300-700 keV energy range) to predict the interaction positions of the scanned super-traces. The performance was compared against PSA using bases derived from the standard simulated ADL bases and simulated bases from AGATAGeFEM.
The position resolution (FWHM) achieved with the three different signal bases for the two known axes of each scan is summarised in Table 1. The experimental NN bases (3.0-3.1 mm) clearly demonstrate superior performance, yielding the best resolution overall. Among the simulated bases, the AGATAGeFEM bases (4.8-6.2 mm) surpassed the performance of the ADL bases(5.2-8.1 mm), a result consistent with previous study [AGATAGeFEM].
@@
Figures 9 and 10 show the PSA error maps for the A005 vertical scan across two different depths: the fourth layer and the back-most sixth layer, respectively. This comparison highlights both consistent performance trends and the impact of interaction depth. Across both layers, several key observations hold true. The direct prediction from the NN model itself (panel (a) in both figures) consistently provides the lowest apparent error and the best homogeneity. Among the PSA-based results, the experimental NN bases (panel b) yield a superior performance compared to the simulated AGATAGeFEM and ADL bases (panels c and d, respectively). The simulated bases, in particular, show degraded performance at radii $\geq$ 20 mm, suggesting potential issues in accurately simulating the core signal rise time in the outer regions of the crystal. A region of increased error is evident in the middle of the segments as it can be seen in layer 4 (Fig. 9 (b)). This error hotspot corresponds to areas where the transient signal strength in neighbouring segments is minimal (see signal at R=22 mm in Fig. 1), reducing the intrinsic position sensitivity. This challenge particularly affects the PSA algorithms, whereas the direct NN model (panel (a)) proves more resilient, exhibiting a significantly smaller increase in error in this region. A direct comparison between the two figures illustrates the general trend of increasing error towards the back of the crystal. The overall error in layer 6 (Fig. 10) is visibly higher for all methods. This is attributed to two main factors: the larger segment volumes at the back reduce position sensitivity, and more importantly, the increased probability of random coincidences and scattering effects from preceding layers in the vertical scan. The performance degradation caused by these effects is significantly more pronounced in the final layer of the crystal.
Figure 9: PSA position error map for the A005 vertical scan across the fourth layer of the crystal. Comparison of the direct NN model prediction (a) and PSA results using NN bases (b), AGATAGeFEM bases (c), and ADL bases (d). Yellow colour is above 8 mm.
Figure 10: PSA position error map for the A005 vertical scan across the sixth (back) layer of the crystal. The panel arrangement is identical to Fig. 9. Yellow colour is above 8 mm.
@@ 3.5 PSA Performance Relative to NN Predictions
Given the potential contamination from random coincidences and mislabelled events in the scanning data, comparing PSA results directly to the scanned position might not be the most accurate way to assess the intrinsic quality of the bases. The NN model itself, trained on the large dataset, might provide a more robust reference position for each super-trace.
Therefore, to further evaluate the performance of each PSA bases, the results were compared to the positions predicted by the NN model itself. The 3D position resolution (FWHM) for each basis is summarised in Table 1.
@@
The analysis of PSA performance is significantly complicated by the presence of mislabelled events, such as random coincidences, which are particularly prevalent towards the back of the crystal in the vertical scan. This is evident when comparing the error maps relative to the scanned position (Fig. 10) with those benchmarked against the NN model’s predictions (Fig. 12). The substantial reduction in error in the latter case demonstrates that comparing PSA results to the NN model’s output effectively neutralises the impact of these random coincidences, as both algorithms agree on the position of the outlier event, independent of the scan label.
@@
Figure 11: Error map comparing PSA predictions to the direct NN model predictions for the A005 vertical scan across the fourth layer. The error is the 2D distance between the position predicted by PSA(bases) and the NN model. Panels show PSA results using: NN bases (a), AGATAGeFEM bases (b), and ADL bases (c). Yellow colour is above 8 mm.
Figure 12: Error map comparing PSA predictions to NN model predictions for the sixth (back) layer. The panel arrangement is identical to the Fig. 11. Yellow colour is above 8 mm.
@@ 4 Discussion
The results presented demonstrate the successful application of neural networks, specifically LSTM-based architectures, for analysing AGATA crystal scanning data obtained with the PSCS scanning method. The key achievement is the development of a method capable of predicting the full 3D interaction position from input super-traces, even when trained using data with only partial (2D) ground truth information available from the vertical and horizontal scans. This was enabled by a custom masked loss function and careful training strategies, including training separate models per segment and incorporating a Gaussian noise layer to mitigate overfitting.
@@
Despite the promising results, certain limitations should be acknowledged, while filtering methods were applied, the training data likely contained residual contamination from random coincidences, multiple-hit events, or other experimental artifacts. The robustness of the NN model against imperfections in the training data was therefore explicitly tested [Abushawish2024]. By deliberately corrupting 50% of the position labels within a segment’s dataset, it was found that the model could still achieve a prediction error close to that obtained with clean data (2.0 mm vs 1.9 mm), particularly when using the masked Euclidean distance loss function, which is inherently less sensitive to outliers than MSE. This result is significant because real experimental scanning data inevitably contains a fraction of mislabelled events. This result increases confidence that the NN model can effectively learn the correct signal-position relationships even from imperfect datasets, supporting the reliability of the generated experimental bases. This resilience mitigates concerns about data purity to some extent.
@@
Comparing the NN approach to the established PSCS algorithm, several advantages emerge. While PSCS provides a valuable algorithmic benchmark, the NN method demonstrated superior signal consistency (lower MSD) and resulted in better PSA performance. Computationally, training the NN models requires significant resources (around 10 hours on 3 NVIDIA RTX6000 GPUs for one crystal), but the subsequent inference step (predicting positions for the entire dataset or generating bases) is relatively fast (approx. 3 hours). For comparison, the standard PSCS algorithm requires roughly 4.7 days to complete the processing on a 2.60 GHz processor machine. Once trained, applying the NN model is straightforward and readily integrated into automated analysis workflows, potentially simplifying the process compared to the potentially complex PSCS algorithm.
@@
An additional useful outcome of applying the NN model is the generation of a large dataset where each experimental super-trace is associated with an estimated 3D interaction position. While the original scanning data only provides partial 2D labels, and simulations might not capture all experimental details, this NN-derived dataset offers comprehensive position estimates for a large portion of the recorded events. This dataset serves two practical purposes. First, it allows for a more detailed characterisation of traditional PSA algorithm performance [Abushawish2024]. By comparing PSA results against the NN model’s 3D position estimates across the detector volume (as explored in Section 3.5), one can gain insights into PSA accuracy and limitations. Second, this dataset provides suitable training material for exploring machine learning models aimed directly at position prediction to replace the current PSA algorithm.
@@
The successful validation using simulated 1D scans opens new possibilities for future characterisation campaigns. If a model can indeed be reliably trained using only one known coordinate per signal, it suggests that scanning could potentially be performed using three orthogonal 1D scans instead of two 2D scans. This could significantly reduce the time required for crystal characterisation. The concept of using 1D scanning for detector characterisation is not new and has been effectively demonstrated in the context of PET scanners using fan-beam collimators [fan_beam, fan_beam1, fan_beam4]. Interestingly, the analysis of that data often relied on machine learning models such as k-nearest neighbour algorithms or gradient-boosted decision trees. This choice of simpler models was well-suited to the nature of PET detector signals, which typically have a smaller size and less complex structure compared to the super-traces from AGATA.
@@ 5 Conclusions
This work successfully demonstrated the development and application of a machine learning approach, utilising LSTM-based neural networks, to analyse experimental data from AGATA crystal scanning campaigns. By employing a masked loss function and tailored training strategies, the NN model effectively learned to predict 3D gamma-ray interaction positions from super-traces despite being trained on data with only 2D ground truth information.
@@
This study highlights the power of machine learning as a tool for processing complex detector signals and overcoming limitations inherent in traditional analysis methods or simulations. The results suggest that NN-based approaches can provide a more accurate and potentially more efficient pathway to generating high-fidelity experimental bases crucial for optimising the performance of position-sensitive detectors like AGATA. Furthermore, the findings provide encouraging, albeit preliminary, evidence for the feasibility of potentially faster 1D scanning techniques for future detector characterisation. This work serves as a foundational step toward developing advanced machine learning-based PSA algorithms capable of resolving multiple interactions within a single segment. Overall, this research underscores the growing role of machine learning in advancing data analysis capabilities in experimental nuclear physics.
@@ Acknowledgments
The authors would like to thank the the Institut Pluridisciplinaire Hubert Curien (IPHC), Strasbourg, for providing the experimental scanning data and the comparative PSCS bases, and the AGATA collaboration for the availability of the S001 and A005 crystals.